# Magnetotransport in the presence of real and momentum space topology

### 13 Mar 2026 v1

@@ I Introduction
Topological semimetals realized in kagome-lattice materials provide a natural platform in which distinct forms of topology can coexist and mutually influence electronic transport [1, 2, 3, 4]. In Weyl semimetals (WSMs), momentum $\mathbf{k}$-space topology arises from isolated band-touching points (Weyl nodes) that act as monopoles of Berry curvature $\bm{\Omega}^{\chi}_{\mathbf{k}}$ with chirality $\chi=\pm 1$ [5, 6]. The associated Berry curvature profoundly modifies semiclassical transport, generating anomalous velocity contributions and phase-space corrections that underlie phenomena such as the anomalous Hall effect and the chiral anomaly [7, 8]. These effects are now well established in Weyl systems [9, 10, 11, 12, 13, 14].
@@
In a kagome WSM hosting both Weyl nodes and a nontrivial magnetic texture, electrons therefore experience two distinct topological fields: the Berry curvature in momentum space and the emergent magnetic field in real space [1, 2, 3, 4]. Within the semiclassical Boltzmann framework, the equations of motion depend only on the total magnetic field $\mathbf{B}_{\mathrm{tot}}=\mathbf{B}+\mathbf{B}_{\mathrm{emer}}$. As a result, the modified phase-space factor that controls the dynamics of the electrons becomes $\mathcal{D}^{\chi}=\left(1+\frac{e}{\hbar}\mathbf{B}_{\mathrm{tot}}\cdot\bm{\Omega}^{\chi}_{\mathbf{k}}\right)^{-1}$[7, 21]. This expression explicitly contains the cross term $\mathbf{B}_{\mathrm{emer}}\cdot\bm{\Omega}^{\chi}_{\mathbf{k}}$, which encodes the interplay between real-space and momentum-space topology. Importantly, the intrinsic Berry curvature of the Weyl nodes remains unchanged; rather, it is the emergent field that modifies the effective magnetic environment in which the Weyl quasiparticles propagate. Consequently, magnetotransport in such systems carries signatures of a mixed real–momentum topological response. In particular, longitudinal magnetoconductivity (LMC) and PHC (PHC) acquire corrections controlled by $\mathbf{B}_{\mathrm{emer}}$, allowing magnetic texture effects to directly influence anomaly-driven transport. Kagome Weyl materials therefore provide an experimentally relevant setting in which this real–momentum topological mixing can be probed through standard magnetotransport measurements. A central conceptual advance of this work is the identification of a mixed topological magnetotransport mechanism arising from the interplay between real-space and momentum-space topology. While previous studies have focused on strain-induced axial fields that are introduced as chirality-dependent gauge fields by construction [22, 14, 23, 24, 25, 26], here the chirality dependence of the transport response emerges dynamically from the coupling between the real-space emergent field and the intrinsic Berry curvature of Weyl fermions.
@@ III Maxwell Boltzmann Transport formalism for electron transport in Skyrmion
$h^{\chi}_{\mathbf{k}}(\theta,\phi)+\frac{\Lambda^{\chi}_{\mathbf{k}}(\theta,\phi)}{\tau_{\mathbf{k}}^{\chi}(\theta,\phi)}$ (17)
@@ IV.1 LMC in the presence of real- and momentum-space topology
Figure 4: Normalized LMC and PHC for a minimal model of untilted WSM as a function of $\gamma$ and $\gamma_{\mathrm{emer}}$. (a) Normalised LMC at $B=0.5$ T. (b) PHC at $B=0.5$ T. (c) Normalized LMC for given $B_{\mathrm{emer}}$ at $B=0$ T. (d) PHC for given $B_{\mathrm{emer}}$ at $B=0$ T. In (a)-(d), from blue to green lines, $B_{emer}$ is varied from 0 to $0.25$ T with $\gamma=\gamma_{\mathrm{emer}}=\pi/2$.
@@ IV.2 PHC in the presence of real- and momentum-space topology
To gain further insight, we consider the PHC in the presence of $B_{\mathrm{emer}}$ alone while varying its direction within the $xz$-plane, as shown in Fig. 4(d). Which shows PHC has very less magnitude, but still finite. It also confirm that $B_{\mathrm{emer}}$ can alone give a finite PHC. It is clear that $\sigma_{zx}\propto\sin 2\gamma_{\mathrm{emer}}$, which is intact for all finite values of $B_{\mathrm{emer}}$ taken. This has never been explored before.
@@ V Conclusions
We have investigated magnetotransport in a time-reversal symmetry–broken, untilted WSM in the simultaneous presence of momentum-space Berry curvature and real-space topology encoded through an emergent magnetic field $\mathbf{B}_{\mathrm{emer}}$. Within a semiclassical Boltzmann framework, we systematically analyzed the LMC and PHC in this mixed-topology setting. In the absence of real-space topology, we recover the known behavior: the LMC undergoes a strong sign reversal as intervalley scattering increases beyond a critical value, reflecting the competition between anomaly-driven charge pumping and relaxation. Upon introducing a finite $B_{\mathrm{emer}}$, the magnetoconductivity acquires an additional shift of its parabolic profile, corresponding to a weak-sign-reversal regime. While intervalley scattering controls the curvature of the response, the emergent field controls the shift of its extremum. Their combined action naturally generates a strong-and-weak sign-reversal regime, demonstrating that real- and momentum-space topology affect distinct geometrical features of the magnetoconductivity.
@@ VI Acknowledgments
The authors are grateful to Amit Agarwal and Gargee Sharma for their insightful discussions and valuable scientific input. A.A. acknowledges financial support from project: SPO/ANRF/PHY/2024434 and IIT Mandi for providing computing facilities. TT acknowledges financial support by the Japan Society for the Promotion of Science, KAKENHI (Grants No. 25H01248) from the Ministry of Education, Culture, Sports, Science, and Technology, Japan.
@@ Appendix A Real and momentum space mixed term in the current
which is linear in $B_{\mathrm{emer}}$ and linear in $B$ (through the standard anomaly correction inside $g^{1}$ when solved self-consistently). Including the usual anomaly contribution $\propto B^{2}$, the conductivity takes the form
$\sigma_{zz}(B)=\sigma_{zz}^{(0)}+a_{1}\left(B+\frac{b_{1}}{2a_{1}}B_{\mathrm{emer}}\right)^{2}-\frac{b_{1}^{2}}{4a_{1}}B_{\mathrm{emer}}^{2}.$ (31)
Hence the extremum shifts to
$B_{0}=-\frac{b_{1}}{2a_{1}}B_{\mathrm{emer}}.$ (32)

### 18 Mar 2026 v2

@@ I Introduction
Topological semimetals realized in kagome-lattice materials provide a natural platform in which distinct forms of topology can coexist and mutually influence electronic transport [1, 2, 3, 4]. In Weyl semimetals (WSMs), momentum $\mathbf{k}$-space topology arises from isolated band-touching points (Weyl nodes) that act as monopoles of Berry curvature $\bm{\Omega}^{\chi}_{\mathbf{k}}$ with chirality $\chi=\pm 1$ [5, 6]. The associated Berry curvature profoundly modifies semiclassical transport, generating anomalous velocity contributions and phase-space corrections that underlie phenomena such as the anomalous Hall effect and the chiral anomaly [7, 8, 9, 10, 11, 12]. These effects are now well established in Weyl systems [13, 14, 15, 16, 17, 18].
@@
In a kagome WSM hosting both Weyl nodes and a nontrivial magnetic texture, electrons therefore experience two distinct topological fields: the Berry curvature in momentum space and the emergent magnetic field in real space [1, 2, 3, 4]. Within the semiclassical Boltzmann framework, the equations of motion depend only on the total magnetic field $\mathbf{B}_{\mathrm{tot}}=\mathbf{B}+\mathbf{B}_{\mathrm{emer}}$. As a result, the modified phase-space factor that controls the dynamics of the electrons becomes $\mathcal{D}^{\chi}=\left(1+\frac{e}{\hbar}\mathbf{B}_{\mathrm{tot}}\cdot\bm{\Omega}^{\chi}_{\mathbf{k}}\right)^{-1}$[7, 25]. This expression explicitly contains the cross term $\mathbf{B}_{\mathrm{emer}}\cdot\bm{\Omega}^{\chi}_{\mathbf{k}}$, which encodes the interplay between real-space and momentum-space topology. Importantly, the intrinsic Berry curvature of the Weyl nodes remains unchanged; rather, it is the emergent field that modifies the effective magnetic environment in which the Weyl quasiparticles propagate. Consequently, magnetotransport in such systems carries signatures of a mixed real–momentum topological response. In particular, longitudinal magnetoconductivity (LMC) and planar Hall conductance (PHC) acquire corrections controlled by $\mathbf{B}_{\mathrm{emer}}$, allowing magnetic texture effects to directly influence anomaly-driven transport. Kagome Weyl materials therefore provide an experimentally relevant setting in which this real–momentum topological mixing can be probed through standard magnetotransport measurements. A central conceptual advance of this work is the identification of a mixed topological magnetotransport mechanism arising from the interplay between real-space and momentum-space topology. While previous studies have focused on strain-induced axial fields that are introduced as chirality-dependent gauge fields by construction [26, 18, 10, 27, 28, 29], here the chirality dependence of the transport response emerges dynamically from the coupling between the real-space emergent field and the intrinsic Berry curvature of Weyl fermions.
@@ III Maxwell Boltzmann Transport formalism for electron transport in Skyrmion
$h^{\chi}_{\mathbf{k}}(\theta,\phi)+\frac{\Lambda^{\chi}_{\mathbf{k}}(\theta,\phi)}{\tau_{\mathbf{k}}^{\chi}(\theta,\phi)}$
@@ IV.1 LMC in the presence of real- and momentum-space topology
Figure 4: Normalized LMC and PHC for a minimal model of untilted WSM as a function of $\gamma$ and $\gamma_{\mathrm{emer}}$. (a) Normalised LMC at $B=0.5$ T. (b) PHC at $B=0.5$ T. (c) Normalized LMC for given $B_{\mathrm{emer}}$ at $B=0$ T. (d) PHC for given $B_{\mathrm{emer}}$ at $B=0$ T. In (a) and (b), from blue to green lines, $B_{emer}$ is varied from 0 to $0.25$ T with $\gamma_{\mathrm{emer}}=\pi/2$. In (c) and (d) $\gamma=\pi/2$
@@ IV.2 PHC in the presence of real- and momentum-space topology
To gain further insight, we consider the PHC in the presence of $B_{\mathrm{emer}}$ alone while varying its direction within the $xz$-plane, as shown in Fig. 4(d). Which shows PHC has very less magnitude, but still finite. It also confirm that $B_{\mathrm{emer}}$ can alone give a finite PHC. It is clear that $\sigma_{zx}\propto\sin 2\gamma_{\mathrm{emer}}$, which is intact for all finite values of $B_{\mathrm{emer}}$ taken. This has never been explored before!
@@ V Conclusions
We have investigated magnetotransport in a time-reversal symmetry-broken, untilted WSM in the simultaneous presence of momentum-space Berry curvature and real-space topology encoded through an emergent magnetic field $\mathbf{B}_{\mathrm{emer}}$. Within a semiclassical Boltzmann framework, we systematically analyzed the LMC and PHC in this mixed-topology setting. In the absence of real-space topology, we recover the known behavior: the LMC undergoes a strong sign reversal as intervalley scattering increases beyond a critical value, reflecting the competition between anomaly-driven charge pumping and relaxation. Upon introducing a finite $B_{\mathrm{emer}}$, the magnetoconductivity acquires an additional shift of its parabolic profile, corresponding to a weak-sign-reversal regime. While intervalley scattering controls the curvature of the response, the emergent field controls the shift of its extremum. Their combined action naturally generates a strong-and-weak sign-reversal regime, demonstrating that real- and momentum-space topology affect distinct geometrical features of the magnetoconductivity.
@@ VI Acknowledgments
The authors are grateful to Amit Agarwal, Gautham Varma K, and Gargee Sharma for their insightful discussions and valuable scientific input. A.A. acknowledges funding from the Core Research Grant by ANRF (Sanction No. CRG/2023/007003), Department of Science and Technology, India. T.T. acknowledges financial support by the Japan Society for the Promotion of Science, KAKENHI (Grants No. 25H01248) from the Ministry of Education, Culture, Sports, Science, and Technology, Japan. The authors are very grateful to IIT Mandi for providing computing facilities.
@@ Appendix A Real and momentum space mixed term in the current
which is linear in $B_{\mathrm{emer}}$ and linear in $B$ (through the standard anomaly correction inside $g^{1}$ when solved self-consistently). Including the usual anomaly contribution $\propto B^{2}$, the conductivity takes the general form
$\sigma_{zz}(B)=\sigma_{zz}^{(0)}+a_{1}B^{2}+b_{1}BB_{\mathrm{emer}},$ (30)
where $a_{1}$ originates from the usual chiral anomaly and $b_{1}$ arises from Eq. (27). Completing the square,
$\sigma_{zz}(B)=\sigma_{zz}^{(0)}+a_{1}\left(B+\frac{b_{1}}{2a_{1}}B_{\mathrm{emer}}\right)^{2}-\frac{b_{1}^{2}}{4a_{1}}B_{\mathrm{emer}}^{2},$ (31)